Previous

Calendar

October

2025

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

28

JD271

29

JD272

30

JD273
Site Visit (HR,MR): HR Details , HR Photos (13)

span cylinder: 1625 psi, N2: 1370 psi

from 10:10-10:58 MST, swapped the 2-micron filter for LI-6262 inlet

from 12:55-1:20 MST, replaced the internal lithium-ion and two 6V batteries in the base of the TC cr23x logger

Turned on NOAH and MRS dataloggers

HR Jeep

bit chilly and breezy. Nice fall day still. scattered clouds

1

JD274

2

JD275

3

JD276

4

JD277

5

JD278

6

JD279

7

JD280
In Cristol (HR)

Recompiled the Edlog lab CR23X programming in a newer version of Edlog to associate array IDs to LoggerNet numerical displays. This works for displaying realtime data in LoggerNet, but requires changes to xml files/something else in the datasystem so the data files can be read properly. Still need to work on the datasystem side before sending the recompiled programming to the CR23X's at the site.

8

JD281

9

JD282
CU "Reading" Day (no classes)

10

JD283

11

JD284

12

JD285

13

JD286

14

JD287

15

JD288
Site Visit (HR, SpB): HR Details SpB Photos (40)

span cylinder: 1550 psi, N2: 1175 psi

from 10:00-12:00 MST, removed REBS Rnet sensor from top of the tower. Moved CNR4 to where the REBS sensor was. Did not change the CNR programming yet.

from 12:30-1:00 MST, added 1 gallon of 50:50 propylene glycol:ethanol, and 26 oz (768 mL) of mineral oil to each pluvio gauge.

brought three new small, span cylinders from LBL to tower (in shed): fb04134, ff66927, fb04006. All at 2000 psi.

Enterprise SUV (HR)

sunny and breezy

16

JD289
at 14:00 MT, Lab Meeting (Nic, MS content)

17

JD290

at 06:52 MST, MET logger stopped sending data to data system

18

JD291

19

JD292

LI-7200 signal strength dropped to 0.1

20

JD293

21

JD294
In lab: (HR, SpB):

Worked on Edlog programming and worked out any kinks in recompiling the program in the data logger. Tested on Lab, MET, and CNR programs and it seems to be working and sending data to the data system, and the data system can read it as well.

at around 11:00 MST, the 21.5m LI-7200 rebooted

22

JD295
Site Visit (HR, AB, HB): HR Details , HR Photos (11)

span cylinder: 1480 psi, N2: 1175 psi

from 10:25-10:33 MST, increased LI-6262 span and N2 delivery pressure so the calibration flow rate matches the sample flow rate.

at 10:40 MST, reconnected the CS I/O cable on the MET logger. At 10:47 MST, sent the recompiled program to the MET logger.

from ~11:20-12:15 MST, replaced the AC/DC converter in Andrew Richardson's CR1000 datalogger box. See notes for more details

added two zip-ties to the turnbuckle connected to the guy wire holding the CNR4 boom up at the top of the tower.

Got NOAH and MRS CR10X's up and running. Resent program to MRS. For NOAH logger, added an AC/DC power supply to power the logger. The battery(ies) seem dead/won't hold a charge.

from ~13:20-14:30 MST, removed LI_7200 sensor head from tower to bring back to the lab. IR source is likely failing.

HR Jeep

sunny and crisp

AmeriFlux Annual Meeting (in Arizona + virtual)

23

JD296
at 14:00 MT, Lab Meeting (Canyon, defense practice)

AmeriFlux Annual Meeting (in Arizona + virtual)

24

JD297
AmeriFlux Annual Meeting (in Arizona + virtual)

25

JD298

26

JD299

27

JD300

28

JD301

29

JD302

30

JD303

31

JD304

1

JD305